OpenStudy (anonymous):

The midpoint of a line segment with end points as (-1, 5) and (x2, 7) is (-4, 6). What is the value of x2? -14 -2 -7 -3

OpenStudy (anonymous):

The midpoint of the segemnt with endpoints (x1, y1) and (x2, y2) is the average for each coordinate. This is the ordered pair:\[(\frac{ x _{1}+x _{2} }{ 2 }, \frac{ y _{1}- y _{2} }{ 2 })\]For the moment, just looking at the y-coordinate, you'll see that the midpoint value of 6 is determined by averaging 7 and 5. So, for the x-coordinate, you average x2 and -1 to get -4. [x2 + (-1)] / 2 = -4 x2 - 1 = -8 so, x2 = -7.

OpenStudy (anonymous):

typo: the ordered pair for y is also supposed to be "+"

OpenStudy (anonymous):

(y1 + y2)/2 that is. Accidentally hit the minus sign.

OpenStudy (anonymous):

thanks Agian Lol Your A Life Saver :D